OpenStudy (anonymous):

Use elimination to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) frac{1}{2}x+y=4 -2x-2y=-6

OpenStudy (anonymous):

Is that supposed to be \[\frac{ 1 }{ 2 }x + y =4?\] If so. We can multiply the top equation by 2 to get \[x+2y=8\] We can then add this to the second equation \[x+2y=8\] +\[-2x-2y=6\] = \[-x = 14\] so x = -14 substitute this back in to get y \[-14+2y=8\] Solve for y to get y = 11. Check to see that we are correct by subbing in: -2(-14)-2(11) = 28-22 = 6. Correct. x = -14, y = 11. As an ordered pair, this is (-14,11)

OpenStudy (anonymous):

awesome! Thank you :)